OpenStudy (anonymous):

I need to solve quotient: 20xy^3/6z divided by 2xy/18z^5 Please help ... I must reduce this quotient to its simplest form.

OpenStudy (anonymous):

\[\frac{ \frac{ 20xy^3 }{ 6z } }{ \frac{ 2xy }{ 18z^5 } }=\frac{ 20xy^3 }{ 6z } *\frac{ 18z^5 }{2xy }\] does that help?

OpenStudy (saifoo.khan):